OpenStudy (anonymous):

step by step 1/3t = 6

OpenStudy (anonymous):

is it\[(1/3)s=6\]or \[1/(3s)=6\]

OpenStudy (saifoo.khan):

Yeah Please Clearly write your questions @hannah!

OpenStudy (anonymous):

sorry its t not s

OpenStudy (anonymous):

there is no s its just MULTIPLY 1/3t=6

OpenStudy (saifoo.khan):

is "t" in the denominator?

OpenStudy (anonymous):

ITS A variable

OpenStudy (saifoo.khan):

I knw, but is it upside or downside?

jhonyy9 (jhonyy9):

1=18t --- t=1/18

OpenStudy (anonymous):

ITS JUST A REGULAR T

OpenStudy (anonymous):

YOU GUYS R CONFUSING MEEE

OpenStudy (anonymous):

(1/3)t=6 3(1/3)t=6*3 t=18 1/(3t)=6 3t=1/6 (1/3)(3)t=(1/6)(1/3) t=1/18 we are asking if the t is in the numerator or the denominator? is it on top of the fraction symbol or on the bottom because u will get different answers

OpenStudy (anonymous):

YOU GUYS R CONFUSING MEEE

OpenStudy (anonymous):

ITS THE TOP ONE

jhonyy9 (jhonyy9):

yes but you have wrote 1/3t and not (1/3)t and this not was clearly sorry